# Game-Theoretic Regularized Self-Play Alignment of Large Language Models

### v1

@@ 4 Regularized Self-Play Policy Optimization
We propose a simple, general, and theoretically sound framework of self-play alignment, namely Regularized Self-Play Policy Optimization (RSPO). Intuitively, RSPO can be understood as an extension of SPPO with an external regularization term, allowing different regularization strategies to be employed. The loss function of RSPO $\mathcal{L}_{\text{RSPO}}$ is defined as the sum of a mean-squared self-play loss, denoted as $\mathcal{L}_{\text{SP}}$, and a weighted regularization term:
@@
where $G({y},\pi_{t},\mu)$, $B(\pi_{t},\mu)$, and $R(\pi_{\theta},\mu)$ are configurable components. First, $G:\mathcal{Y}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ defines the update direction of $\pi_{\theta}$, which can be set as the gradient of a utility function to guide the policy update towards increasing the utility. Second, the baseline function $B:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ is for variance-reduction for $G$, similar to the baseline in REINFORCE (Williams, 1992). Lastly, $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is the regularization function. The coefficient $\lambda$ is the regularization temperature. The first Mean Square Error term in Equation 7 can be interpreted as a self-play loss of conducting exponentiated gradient descent (Beck and Teboulle, 2003).
RSPO offers a simple way to introduce regularization into self-play alignment with only an additional term in the loss. Using RSPO, self-play can be regularized by simply adding a term $R$ to the self-play loss functions. For example, to regularize SPPO (Equation 4), we can leverage RSPO loss function (Equation 7) by setting appropriate $G$ and $B$, and directly adding an additional term of $R$ to the loss function of SPPO. Therefore, RSPO offers the simplicity and flexibility to incorporate various regularization methods into self-play-based preference optimization methods.
Furthermore, RSPO is a generalized and versatile framework that makes the import of regularization into existing self-play methods more efficient. We show in Section 4.1 that RSPO can generalize existing self-play methods with $\mathcal{L}_{\text{SP}}$ and without external regularization $R$. Thus, regularizing existing methods requires no change to their loss functions or hyperparameters, but simply adding an external regularization to their loss function and tuning the temperature $\lambda$.
We provide the theoretical guarantee for RSPO in Section 4.2 by showing that RSPO is the RL implementation of a specific type of Mirror Descent. Given a specific regularizer $R$, RSPO has a last-iterate convergence to the NE of the corresponding regularized game. We also introduce the implementation details of RSPO specifically for preference optimization (Equation 5) in Section 4.3.
@@ 4.1 Generalizing Existing Self-Play Methods
We show that existing methods have loss functions equivalent to RSPO without external regularization: $\mathcal{L}_{\text{RSPO}}(\theta;G,B,R=0)$. First of all, unregularized self-paly method SPPO (Wu et al., 2024), has loss function in Equation 4 satisfying:
$\mathcal{L}_{\text{SPPO}}(\theta)=\mathcal{L}_{\text{RSPO}}\bigg{(}\theta;G=\mathbb{P}(y\succ\pi_{t}),B=\frac{1}{2},R=0\bigg{)}.$ (8)
Other unregularized self-play methods following the preference-based exponential update in Equation 3 can also be generalized by $\mathcal{L}_{\text{SP}}$, and thus can be regularized by simply adding regularization term to the loss functions. Based on the same exponential update rule as in SPPO, SPO (Swamy et al., 2024) is equivalent to updating policy with the loss in Equation 8. Magnetic Policy Optimization (Wang et al., 2024), despite incorporating regularization in the policy update, periodically update $\mu=\pi_{t}$. Consequently, it inherently follows update in Equation 3 while incorporating multiple policy updates within each iteration, following the approach of (Tomar et al., 2020).
In addition, even existing regularized methods can be generalized by $\mathcal{L}_{\text{RSPO}}$ without external regularization. Mirror Descent methods including Online Mirror Descent and Nash-MD (implemented with RL) have direct connection to RSPO (the detailed derivations are provided in Appendix A.1):
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)=\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\bigg{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu})-\tau\log\frac{\pi_{t}(y)}{\mu(y)},B=\frac{1}{2},R=0\bigg{)}.$ (9)
This equivalence fundamentally arises from the well-established connection between the reinforcement learning (RL) policy gradient and the gradient of a quadratic loss function, as explored in various works (Haarnoja et al., 2018; Tomar et al., 2020; Malkin et al., 2022). We summarize the generalization of RSPO in Table 3. Therefore, our generalized loss framework RSPO enables to even add extra regularization to existing regularized self-play methods.
4.2 Theoretical Results
In this section, we examine the theoretical properties of RSPO, with a particular emphasis on its convergence guarantee. We adopt Mirror Descent (MD) as the foundational framework, given its well-established last-iterate convergence to the NE in game-solving.
@@ 4.2 Theoretical Results
Notably, the vanilla Magnetic Mirror Descent limits $R$ to be the same regularization method of $\pi$ and $\pi_{t}$, i.e., $R=B_{\psi}$ (Sokota et al., 2022, Section 3.2); whereas in this paper we aim at a general regularizer of $\pi$ and $\mu$, which could be different from $B_{\psi}$, and study the effects of different regularizations methods.
@@
If $R(\cdot,\mu)$ is $1$-strongly convex relative to $\psi$, $\eta\leq\tau$, and $U$ is linear, then policy updated by GMMD in Equation 10 has last-iterate convergence to the following regularized NE:
Proposition 4.1 is a direct application of Theorem 3.4 by Sokota et al. (2022), which guarantees the last-iterate convergence of GMMD to the NE of a regularized game. We provide the proof in Appendix A.3.
Deep RL Implementation of GMMD. To adapt GMMD to preference optimization problems, RL techniques are commonly employed as practical implementations, as for many MD update (Tomar et al., 2020; Munos et al., 2023; Wang et al., 2024). Define the loss function of conducting GMMD in preference optimization as
@@
Here, we set the Bregman divergence to Reverse KL in preference optimization as in previous works (Munos et al., 2023; Zhang et al., 2024). The gradient estimation of $\mathcal{L}_{\text{GMMD}}(\theta)$ for policy updates is required since the expectation in the first term is dependent on $\pi_{\theta}$. Following Policy Gradient theorem (Sutton et al., 1999), then we have
@@
We follow SPPO to replace the samples $y\sim\pi_{\theta}$ with $y\sim\pi_{t}$ directly since they are equivalent while computing the loss before updating, and rewrite the loss equivalent to GMMD:
$\nabla_{\theta}\mathcal{L}_{\text{GMMD}}(\theta)=\nabla_{\theta}\Bigg{(}\frac{1}{2}{\mathbb{E}_{{y\sim\pi_{t}}}\Big{[}-\eta G(y;\pi_{t})+\log\frac{\pi_{\theta}(y)}{\pi_{t}(y)}{+\eta B}\Big{]}^{2}}+\tau R(\pi_{\theta},\mu)\Bigg{)}.$ (14)
Therefore, according to Equation 14, RSPO is equivalent to the RL implementation of GMMD since the RHS of the Equation 14 is RSPO loss up to multiplying a constant.
4.3 Implementation for Preference Optimization
In this section, we introduce the implementation of RSPO. We set the update direction $G$ to be the gradient of the preference against $\pi_{t}$, $\forall y\in\mathcal{Y}$:
As for baseline function $B$, we set $B=\tfrac{1}{2}$ following Nash-MD and SPPO. In theory, $B$ helps minimize the variance of $G$ the most when $B=\mathbb{E}_{y\sim\pi_{t}}[G(y,\pi_{t},\mu)]$. But in preference optimization, due to the typically small minibatch size, the estimation error of the mean of $G$ could be large, leading to additional estimation error of the loss. Thus, we also set the baseline value for variance reduction to be a constant $\tfrac{1}{2}$, the mean value of $G$ when the algorithm is converged.
Specifically, we execute Algorithm 1 by applying the following RSPO loss with any regularization $R$ of interest:
$\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}),B=\tfrac{1}{2},R\big{)}=\mathcal{L}_{\text{SPPO}}+\lambda R(\pi_{\theta},\mu).$ (16)
Therefore, we can directly use the off-the-shelf self-play alignment method SPPO to implement RSPO. The only modification required is to add the regularization to the SPPO loss. Various divergences for regularization are implemented, as discussed in Appendix B.3.
For completeness, we provide the convergence guarantee for this instance of RSPO to the Nash equilibrium of the regularized preference optimization game as follows (Proof in Appendix A.4).
@@ Methods
Mistral-7B-Instruct (Jiang et al., 2023a) 17.1 12.6 7.51
@@
Table 1: Performance of existing methods, and our strongest model: RSPO with Importance-Sampling-based Forward KL ($\lambda_{1}=0.1$) + Reverse KL ($\lambda_{2}=0.5$) divergence as regularization, on AlpacaEval-2 and Arena-Hard-Auto-v0.1.
  • Q1: Does regularization improve the performance of self-play alignment of LLMs (Sec. 5.1)?
  • Q3: What additional advantages can be derived from utilizing regularization in self-play (Sec. 5.3)?
Experiment Setup. We investigate our methods mainly on benchmarks AlpacaEval (Dubois et al., 2024), Arena-Hard (Li et al., 2024), and MT-Bench (Zheng et al., 2023). We follow the experiment setup of SPPO and Snorkel-Mistral-PairRM-DPO (Snorkel) (Tran et al., 2023) to examine our regularization methods, where Snorkel is based on iterative DPO and has achieved strong performance on AlpacaEval. Our reference policy model is Mistral-7B-Instruct-v0.2. Since iterative self-play methods require no response data for training, we only use the prompts of the Ultrafeedback dataset (Cui et al., 2023), whose size is $\sim 60$K. Following SPPO and Snorkel, we split the prompts into three subsets and use only one subset per iteration to prevent over-fitting. To understand the later-iterate performance of self-play, in section 5.1, we also train on the single fold of the prompts iteratively. We use a 0.4B response-pair-wise preference model PairRM (Jiang et al., 2023b), evaluated as comparable to 10×10$\times$10 × larger reward/preference models (Cui et al., 2023). We investigate the effect of regularization mainly via AlpacaEval-2.0, where the main metric is length-controlled win rate (LCWR).
Implementations and Baselines. The implementation of self-play methods follows Algorithm 1. In each iteration, given response-pair-wise preference from PairRM and $K=5$ number of response samples from the current policy, we estimate the policies’ preference $\mathbb{P}(\pi\succ\pi_{t})$ and regularization via Monte-Carlo estimation to compute the loss function. We replicate the SPPO with the default hyper-parameters and extend it to $9$ iterations. We implement RSPO as described in Corollary 4.2. The implementation of regularizations in RSPO is demonstrated in Appendix B.3 using the $K$ samples. We report some of the baseline results from the previous papers, including SPPO, Snorkel (Mistral-PairRM-DPO) (Tran et al., 2023), Mistral-7B (Instruct-v0.2) (Jiang et al., 2023a), iterative DPO by Wu et al. (2024), and SimPO Meng et al. (2024). Since the SPPO paper only provides results across $3$ iterations (Wu et al., 2024), we replicate SPPO as an important baseline to study the performance across more than $3$ iterations.
5.1 Effectiveness of Regularization
Figure 2: Left: LC win rate across iterations for unregularized self-play method SPPO, SPPO trained on a subset of the data (SPPO (subset)), and reverse-KL-regularized RSPO (RSPO (Rev. KL)). The base model is Mistrial-7B. SPPO starts to degrade after $3$ iterations. Right: LC win rate of SPPO and RSPO with different regularization methods. From left to right regularization methods: Reverse KL ($\lambda=0.5$), Forward KL ($\lambda=1.0$), Chi-Squared ($\lambda=0.1$), Importance-Sampling Forward KL ($\lambda=0.1$), Forward and Reverse KL linear combination ($\lambda_{1}=0.1$, $\lambda_{2}=0.5$). RSPO outperforms unregularized self-play (SPPO) significantly.
In this section, we assess the effectiveness of regularization primarily by comparing the performance of unregularized and regularized self-play methods. We first examine the over-optimization issue inherent in practical self-play preference optimization by extending the execution of SPPO to Iteration 5. As depicted in Figure 2, a decline in performance appears during the later iterations of SPPO. We hypothesize that this behavior arises from the practical challenges associated with a misspecified preference model, as the signals driving policy updates in SPPO rely only on the preference model.
In Table 2, we further contrast the unregularized self-play method, SPPO, and other iterative methods with the best RSPO, namely RSPO (For.+Rev.). The regularization is a linear combination of Forward KL and Reverse KL divergence with coefficients $0.1$ and $0.5$, respectively. The comparative results reveal that regularization enhances the SPPO win rate from $31.02\%$ to $38.31\%$, and the LC win rate increases from $28.53\%$ to $35.44\%$ in iteration $3$. Notably, in the first iteration, reg. SPPO exhibits a slightly lower LC win rate, potentially attributable to the influence of strong regularization. However, subsequent iterations show a marked improvement, with the LC win rate of reg. SPPO increases by up to $7.53\%$ within a single iteration. In summary, the findings in Table 2 underscore the effectiveness of regularization in self-play optimization.
Additionally, to rule out the possibility of insufficient iterations affecting performance, we report the best result among nine iterations of our replicated SPPO in Table 2, denoted as "SPPO $\leq 9$". SPPO $\leq 9$ consistently underperforms the RSPO result at iteration $3$. These observations emphasize that even extended training under the unregularized framework fails to match the performance gains achieved through regularization, thereby affirming the critical role of regularization in self-play methodologies for preference optimization.
@@ 5.1 Effectiveness of Regularization
Regularization Iteration AlpacaEval 2.0

### v2

@@ 3 Regularized Self-Play Policy Optimization
We propose a framework for regularized self-play alignment, namely Regularized Self-Play Policy Optimization (RSPO). RSPO is simple and flexible for regularization, and provably convergent to Nash Equilibrium. The loss function of RSPO $\mathcal{L}_{\text{RSPO}}$ is defined as the sum of a mean-squared self-play loss and a weighted regularization term:
@@
where $G({y},\pi_{t},\mu)$, $B(\pi_{t},\mu)$, and $R(\pi_{\theta},\mu)$ are configurable components. First, $G:\mathcal{Y}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ defines the update direction of $\pi_{\theta}$, which can be set as the gradient of a utility function to guide the policy update towards increasing the utility. Second, the baseline function $B:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow(-\infty,\infty)$ is for variance-reduction of $G$, similar to the baseline in REINFORCE (Williams, 1992). Lastly, $R:\Delta^{\mathcal{X}}_{\mathcal{Y}}\times\Delta^{\mathcal{X}}_{\mathcal{Y}}\rightarrow\mathbb{R}$ is the regularization function. The coefficient $\lambda$ is the regularization temperature. The first Mean Square Error term in Equation 6 can be interpreted as a self-play loss of conducting exponentiated gradient descent (Beck and Teboulle, 2003).
RSPO is a modular framework offering a simple way to introduce regularization into self-play alignment with only an additional term in the loss. RSPO offers the simplicity and flexibility to incorporate various regularization methods into self-play-based preference optimization methods. Additionally, we show in Section 3.1 that RSPO can generalize existing unregularized self-play methods without external regularization $R$. Thus, regularizing existing methods requires no change to their original loss functions or hyperparameters, but simply adding an external plug-and-play regularization to their loss function and tuning the temperature $\lambda$.
In practice, we set baseline function $B=\tfrac{1}{2}$ following Nash-MD and SPPO, and the update direction $G$ to be the gradient of the preference against $\pi_{t}$, $\forall y\in\mathcal{Y}$:
We execute Algorithm 1 by applying the following RSPO loss with any regularization $R$ of interests:
$\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}),B=\tfrac{1}{2},R\big{)}.$ (8)
In theory, $B$ helps minimize the variance of $G$ the most when $B=\mathbb{E}_{y\sim\pi_{t}}[G(y,\pi_{t},\mu)]$. But in preference optimization, due to the typically small minibatch size, the estimation error of the mean of $G$ could be large, leading to additional estimation error of the loss. Thus, we also set the baseline value for variance reduction to be a constant $\tfrac{1}{2}$, the mean value of $G$ when the algorithm converged. For the implementation of various divergence-based regularization, refer to Appendix C.3.
In the following sections, we first illustrate the generalizable formulation of RSPO, so that it can be implemented without modifying the existing self-play component. We then establish theoretical convergence guarantees for RSPO grounded in Mirror Descent theory.
@@ 3.1 Generalizing Existing Self-Play Methods
In this section, we show how RSPO generalize existing self-play methods, which showcase (1) implementing RSPO requires only one additional term to existing self-play loss functions; (2) the limitation of existing regularized methods. First, the unregularized self-play method SPPO (Wu et al., 2024) has a loss function defined in Equation 3 equivalent to RSPO without external regularization:
[…]
According to Equation 8 and Equation 9, $\mathcal{L}_{\text{RSPO}}=\mathcal{L}_{\text{SPPO}}+\lambda R(\pi_{\theta},\mu)$, i.e. the implementation of RSPO is equivalent to directly add the regularization $R$ to the loss function of SPPO (Equation 3). This implies that the additional regularization term becomes plug-and-play, requiring minimal changes to existing training pipeline.
In addition, existing regularized methods can be generalized by $\mathcal{L}_{\text{RSPO}}$ (derivations in Appendix A.2):
$\nabla_{\theta}\mathcal{L}_{\text{Nash-MD}}(\theta)=$ $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu}),B=\frac{1}{2},R=D_{\text{KL}}(\pi_{\theta}\|\mu)\big{)}$ (10)
$=$ $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}\Big{(}\theta;G=\mathbb{P}(y\succ\pi_{t}^{\mu})-\tau\log\frac{\pi_{t}(y)}{\mu(y)},B=\frac{1}{2},R=0\Big{)}.$ (11)
Equation 10 verifies our summarization shown in Figure 1. The convergence guarantee of Nash-MD (Munos et al., 2023, Lemma 2) requires the policy updated with Equation 5, which is specifically designed for reverse KL regularization, as other $R$ can not be merged with $D_{\text{KL}}(\pi\|\mu)$ to a regularization w.r.t. geometric mixture $\pi^{\mu}_{t}$. Additionally, Equation 11 demonstrates that RSPO enables to even add extra regularization to existing regularized self-play methods.
3.2 Theoretical Guarantees
In this section, we examine the theoretical properties of RSPO, with a particular emphasis on its convergence guarantee. We adopt Mirror Descent (MD) as the foundational framework, given its well-established last-iterate convergence to the NE.
@@ 3.2 Theoretical Guarantees
Tabular GMMD. Denote the utility function of the game as $U$, define $G$ as the element of the vector of partial derivatives of $U$ w.r.t. policy:
@@
Notably, the vanilla Magnetic Mirror Descent limits $R$ to be the same regularization method of $\pi$ and $\pi_{t}$, i.e., $R=B_{\psi}$ (Sokota et al., 2022, Section 3.2); whereas in this paper we aim at a general regularizer of $\pi$ and $\mu$, which could be different from $B_{\psi}$, and study the effects of different regularization methods.
@@
If $R(\cdot,\mu)$ is $1$-strongly convex relative to $\psi$, $\eta\leq\tau$, and $U$ is linear, then policy updated by GMMD in Equation 13 has last-iterate convergence to the following regularized NE $\max_{\pi}\min_{\pi^{\prime}}U(\pi;\pi^{\prime})-\tau R(\pi,\mu)+\tau R(\pi^{\prime},\mu).$
Proposition 3.1 is a direct application of Theorem 3.4 by Sokota et al. (2022), which guarantees the last-iterate convergence of GMMD to the NE of a regularized game (Proof in Appendix A.3).
Deep RL Implementation of GMMD. To adapt GMMD to preference optimization problems, RL techniques are commonly employed as practical implementations, as for many MD update (Tomar et al., 2020; Munos et al., 2023; Wang et al., 2024a). Define the loss function of conducting GMMD in preference optimization as
@@
Here, we set the Bregman divergence to Reverse KL in preference optimization as in previous works (Munos et al., 2023; Zhang et al., 2024). The gradient estimation of $\mathcal{L}_{\text{GMMD}}(\theta)$ for policy updates is required since the expectation in the first term is dependent on $\pi_{\theta}$. Following Policy Gradient (PG) theorem (Sutton et al., 1999), the PG of GMMD is equal to $\nabla_{\theta}\mathcal{L}_{\text{RSPO}}(\theta)$ up to multiplying a constant:
@@
Due to the equivalence between RSPO and GMMD, we provide the convergence guarantee for our practical implementation of RSPO (Equation 13), to the Nash equilibrium of the regularized preference optimization game as follows (Proof in Appendix A.4).
@@ 4 Experiments
  • Q1: Does regularization improve the performance of self-play alignment (Sec. 4.1)?
  • Q3: What additional advantages can be obtained by regularization in self-play (Sec. 4.3)?
Experiment Setup. We evaluate our methods mainly on benchmarks AlpacaEval (Dubois et al., 2024), Arena-Hard (Li et al., 2024), and MT-Bench (Zheng et al., 2023), and test the response generation diversity and quality via self-BLEU (Zhu et al., 2018) and ArmoRM (Wang et al., 2024b), respectively. We follow the experiment setup of SPPO and Snorkel-Mistral-PairRM-DPO (Snorkel) (Tran et al., 2023) to examine our regularization methods, where Snorkel is based on iterative DPO and has achieved strong performance on AlpacaEval. Our reference policy model and base model are both Mistral-7B-Instruct-v0.2. Since iterative self-play methods require no response data for training, we only use the prompts of the Ultrafeedback dataset (Cui et al., 2023), whose size is $\sim 60$K. Following SPPO and Snorkel, we split the prompts into three subsets and use only one subset per iteration to prevent over-fitting. To understand the later-iterate performance of self-play, in section 4.1, we also train on the single fold of the prompts iteratively. We use a 0.4B response-pair-wise preference model PairRM (Jiang et al., 2023a), evaluated as comparable to 10×10$\times$10 × larger reward/preference models (Cui et al., 2023).
Implementations and Baselines. The implementation of self-play methods follows Algorithm 1. In each iteration, given response-pair-wise preference from PairRM and $K=5$ number of response samples from the current policy, we estimate the policies’ preference $\mathbb{P}(\pi\succ\pi_{t})$ and regularization via Monte-Carlo estimation to compute the loss function. We replicate SPPO with the default hyper-parameters and extend it to $9$ iterations. We implement RSPO as described in Corollary 3.2. The implementation of regularizations in RSPO is demonstrated in Appendix C.3 using the $K$ samples. We report some of the baseline results from the previous papers, including SPPO, Snorkel (Mistral-PairRM-DPO) (Tran et al., 2023), Mistral-7B (Instruct-v0.2) (Jiang et al., 2023b), iterative DPO by Wu et al. (2024), and SimPO Meng et al. (2024). Since the SPPO paper only provides results across $3$ iterations (Wu et al., 2024), we replicate SPPO as an important baseline to study the performance across more than $3$ iterations.
Figure 2: Left: Length-controlled win rate (LCWR) across iterations for unregularized self-play method: SPPO, SPPO trained on a subset of the data: SPPO (subset), and reverse-KL regularized RSPO: RSPO (Rev. KL). SPPO starts to degrade from iteration 3 due to overoptimization, RSPO with reverse KL regularization mitigates it. Right: LCWR of SPPO and RSPOs in AlpacaEval-2.0. RSPO (IS-For.+Rev.) outperforms SPPO and others including reverse KL regularization.
4.1 Effectiveness of Regularization
In this section, we assess the effectiveness of regularization primarily by comparing the performance of unregularized and regularized self-play methods. We first examine the over-optimization issue inherent in practical self-play alignment by extending the execution of SPPO to Iteration 5. As depicted in Figure 2 (left), a decline in performance appears during the later iterations of SPPO. We hypothesize that this behavior arises from the practical challenges associated with over-optimization.
We then present comprehensive results across three widely used benchmarks (Table 1). RSPO with forward and reverse KL regularization, consistently outperforms the unregularized baseline (SPPO) [^2] We report our replicated testing of the published SPPO Iter3 model (link) on Arena-Hard benchmark. Thus, it is different from the result presented in the original paper of SPPO (Wu et al., 2024)., and iterative DPO in iteration 3, and the strong offline method SimPO across all benchmarks, with a clear performance margin. These results underscore the importance of incorporating regularization into self-play alignment. We hypothesize that the effectiveness of regularization arises from the continued utility of the reference policy during optimization, which provides stable guidance and helps mitigate inaccuracies or misspecifications in the general preference model (PairRM).
@@ Methods
Mistral-7B-Instruct (Jiang et al., 2023b) 17.1 12.6 7.51
@@
Table 1: Comprehensive comparisons on three popular benchmarks of baselines, and our strongest model. RSPO with Importance-Sampling-based Forward KL ($\lambda_{1}=0.1$) + Reverse KL ($\lambda_{2}=0.5$) divergence as regularization outperforms baselines on all benchmarks with a clear margin.
@@
Table 2: AlpacaEval LCWR of iterative methods. RSPO shows fast improvement over iterations.
In Table 2, we further contrast the performance dynamics across iterations of methods: unregularized self-play method, (SPPO), other iterative methods, and the best RSPO, namely RSPO (For.+Rev.), regularized by the linear combination of Forward KL and Reverse KL divergence with temperatures of $0.1$ and $0.5$, respectively. The comparative results reveal that regularization enhances the SPPO win rate from $31.02\%$ to $38.31\%$, and the LC win rate increases from $28.53\%$ to $35.44\%$ in iteration $3$. Notably, in the first iteration, reg. SPPO exhibits a slightly lower LC win rate, potentially attributable to the influence of strong regularization. However, subsequent iterations show a marked improvement, with the LC win rate of reg. SPPO increases by up to $7.53\%$ within a single iteration. In summary, Table 2 underscores the effectiveness of regularization in self-play optimization.
[…]
4.2 Impact of Different Regularizations
We then study the effect of applying different regularization $R$ in RSPO. To obtain a well-regularized self-play, the tuning of regularization temperature $\lambda$ is necessary. An ablation study of the regularization temperature of different methods is shown in Figure 3. According to the figure, the response length increases along with the temperature in reverse KL divergence and Chi-square divergence regularized RSPO. However, both the length and win rate are decreased with stronger regularization via Forward KL divergence, implemented using importance sampling. We attribute the decreasing win rate to the violation of relative convexity assumption (A.1), and the length reduction to the intrinsic mass-averaging property of forward-KL divergence divergence when used for regularization.
In particular, the raw win rate analysis highlights reverse KL divergence as a crucial factor in enhancing self-play performance. We attribute the observed effect to the inherent mode-seeking behavior of reverse KL divergence. Given that forward KL divergence tends to reduce response length while reverse KL divergence yields significant improvements, we adopt a linear combination of both. This approach is designed to balance their complementary effects, ultimately optimizing for a higher LCWR (RSPO (IS-For. + Rev.) in Figure 2 Right). The hyperparameters provided in Table 6.
Figure 3: Ablation study on regularization temperature $\lambda$ of RSPO conducted on AlpacaEval 2.0 prompts. We evaluate how the average response length and raw WR are affected by the regularization temperature. Higher temperature of forward KL leads to shorter response length.
4.3 Response Diversity and Other Aspects
Regularization Iteration AlpacaEval 2.0 Dataset